### v1

@@ 7 Impulse and Singular Control Connections
    $F(K):=F(K;w,y)=\bigg(\frac{pB\mathfrak{i}}{B\xi}+\frac{\gamma Bg}{B\xi}-\frac{K}{B\xi}\bigg)(w,y),\hskip 18.49988pt(w,y)\in\mathcal{R}.$ (7.1)
@@
$h(y_{K}^{*})=F(w_{K}^{*},y_{K}^{*})=\frac{\gamma Bg(w_{K}^{*},y_{K}^{*})+pB\mathfrak{i}(w_{K}^{*},y_{K}^{*})-K}{B\xi(w_{K}^{*},y_{K}^{*})}=\frac{\gamma g^{\prime}(\theta_{1})+p}{\xi^{\prime}(\theta_{1})}-\frac{K}{B\xi(w_{K}^{*},y_{K}^{*})}$
@@ 8 Conclusions and Remarks
This paper presents a comprehensive analysis of a class of long-term average impulse control problems having two sources of income, together with their closely related singular control formulations. Working within a general one-dimensional diffusion framework with boundary conditions motivated by applications, we derive explicit solutions and characterized the optimal controls for both problems. Our findings rigorously establish an intrinsic connection between the two control formulations, with the singular control solution emerging as a natural limit of the impulse control problem as fixed costs vanish. It also identifies the penalty incurred on the optimal impulse value (relative to the optimal singular value) by the presence of the positive fixed cost $K$. This penalty includes a reduction due to the restriction to the subclass of impulse controls along with the long-term average reduction arising directly from the fixed costs.
@@ A.2 Proofs of Section 3
$\lim_{x\to b}f(x)$ $=\lim_{x\to b}(\xi(x)-\xi(w))\bigg[z-\frac{x-w}{\xi(x)-\xi(w)}\bigg]=\infty.$ (A.8)
@@
$=\limsup_{w\to a}\bigg(\frac{g(w_{\varepsilon})-g(w)}{\xi(w_{\varepsilon})-\xi(w)}+\frac{g(y_{0})-g(w_{\varepsilon})}{\xi(w_{\varepsilon})-\xi(w)}\bigg)\cdot\frac{1}{1+\frac{\xi(y_{0})-\xi(w_{\varepsilon})}{\xi(w_{\varepsilon})-\xi(w)}}$

### v2

@@ 7 Impulse and Singular Control Connections
    $F(K):=F(K;w,y)=\left(\frac{pB\mathfrak{i}}{B\xi}+\frac{\gamma Bg}{B\xi}-\frac{K}{B\xi}\right)(w,y),\hskip 18.49988pt(w,y)\in\mathcal{R}.$ (7.1)
@@
$h(y_{K}^{*})$ $=F(w_{K}^{*},y_{K}^{*})=\frac{\gamma Bg(w_{K}^{*},y_{K}^{*})+pB\mathfrak{i}(w_{K}^{*},y_{K}^{*})-K}{B\xi(w_{K}^{*},y_{K}^{*})}$
$=\frac{\gamma g^{\prime}(\theta_{1})+p}{\xi^{\prime}(\theta_{1})}-\frac{K}{B\xi(w_{K}^{*},y_{K}^{*})}=h(\theta_{1})-\frac{K}{B\xi(w_{K}^{*},y_{K}^{*})}$
@@ 8 Conclusions and Remarks
This paper presents a comprehensive analysis of a class of long-term average impulse control problems having two sources of income, together with their closely related singular control formulations. Working within a general one-dimensional diffusion framework with boundary conditions motivated by applications, we derive explicit solutions and characterize the optimal controls for both problems. Our findings rigorously establish an intrinsic connection between the two control formulations, with the singular control solution emerging as a natural limit of the impulse control problem as fixed costs vanish. It also identifies the penalty incurred on the optimal impulse value (relative to the optimal singular value) by the presence of the positive fixed cost $K$. This penalty includes a reduction due to the restriction to the subclass of impulse controls along with the long-term average reduction arising directly from the fixed costs.
@@ A.2 Proofs of Section 3
$\lim_{x\to b}f(x)$ $=\lim_{x\to b}(\xi(x)-\xi(w))\left[z-\frac{x-w}{\xi(x)-\xi(w)}\right]=\infty.$ (A.8)
@@
$=\limsup_{w\to a}\left(\frac{g(w_{\varepsilon})-g(w)}{\xi(w_{\varepsilon})-\xi(w)}+\frac{g(y_{0})-g(w_{\varepsilon})}{\xi(w_{\varepsilon})-\xi(w)}\right)\cdot\frac{1}{1+\frac{\xi(y_{0})-\xi(w_{\varepsilon})}{\xi(w_{\varepsilon})-\xi(w)}}$